OpenStudy (anonymous):

What is the 32nd term of the arithmetic sequence where a1 = 15 and a13= -57?

OpenStudy (amistre64):

if we shift the index to start at a0 ... this might be easier to see

OpenStudy (amistre64):

a0 = 15 +d(0) a1 = 15 + d(1) a2 = 15 + d(2) a3 = 15 + d(3) a4 = 15 + d(4) a5 = 15 + d(5) an = 15 + dn ; but keep in mind the shift

OpenStudy (amistre64):

a13 is a12: a12 = -57 -57 = 15 + d(12) solve for d

OpenStudy (amistre64):

then use it to determine: a32 is shifted to a31 a31 = 15 + d(31)

OpenStudy (anonymous):

d=-6

OpenStudy (anonymous):

a31=-171

OpenStudy (anonymous):

Thank you!

OpenStudy (amistre64):

correct :)